Which circles have an area less than 300 square centimeters?
1
Expert's answer
2020-05-05T20:11:45-0400

Let's write a formula for circle area:

A = "\pi*R^2"

To answer this question we just need to find the radius:

R = "\sqrt{\dfrac{A}{\pi}}=\sqrt{\dfrac{300}{3.14}}\approx9.77cm"

So, all circles that have radius between 0 and 9.77cm will have area less than 300cm2

Or you can write it like: 0 < R < 9.77